can you help me and thankn you in advanced I just can't ge t it
find x-andy- intercept of f(x)=(x+4)(2x-1)(x-1)and last can you
write the equation of the line with a slope slope of 2/3 and y-intercept of 1( write in intercept form. thank you again

Answer by stanbon(57361)

find x-and y- intercept of f(x)=(x+4)(2x-1)(x-1)
-----------
x-intercept:
Let y = 0 and solve for "x":
(x+4)(2x-1)(x-1) = 0
x = 4 or 1/2 or 1
Those are the x-intercepts.
------------------------------------
y-intercept:
Let x = 0 and solve for "y":
y = (0+4)(0-1)(0-1) = 4
that is the y-intercept
--------------------------------------
and last can you
write the equation of the line with a slope slope of 2/3 and y-intercept of 1( write in intercept form.
---
Slope = 2/3
intercept = 1
Equation:
y = (2/3)x + 1
=======================
Cheers,
Stan H.